OpenStudy (anonymous):

Can someone please help me with this question? Please and Thank you. Solve the equation v + f - e = 2 for e. A. e = v - f - 2 B. e = v + f - 2 C. e = -v - f + 2 D. e = -v + f + 2

OpenStudy (souvik):

first subtract f and v from both sides

OpenStudy (texaschic101):

do you understand what souvik is saying to do ?

OpenStudy (anonymous):

I'm lost?

OpenStudy (anonymous):

Nope. @texaschic101

OpenStudy (texaschic101):

v + f - e = 2 for e first we will subtract v from both sides v - v + f - e = 2 - v combine like terms f - e = 2 - v are you with me so far ?

OpenStudy (texaschic101):

we are trying to get e on one side and everything else on the other side

OpenStudy (anonymous):

yes. @texaschic101

OpenStudy (texaschic101):

f - e = 2 - v now we will subtract f from both sides f - f - e = 2 - v - f combine like terms -e = 2 - v - f with me so far ?

OpenStudy (anonymous):

Is it B? e = v + f - 2?

OpenStudy (souvik):

yeah! right..how did u do that?@mercedesjaay_

OpenStudy (texaschic101):

-e = 2 - v - f multiply the equation by -1 to make e positive (-1)-e = 2 - v - f e = v + f - 2 you are correct :)

OpenStudy (anonymous):

I followed @texaschic101 Thanks a bunch! :)

OpenStudy (texaschic101):

anytime :)